# ExCyTIn-Bench: Evaluating LLM agents on Cyber Threat Investigation

### v1

@@ C.1 Features
Real-world security database Although the tenant is for a fictional company, the log data we use is from real-world tenant environments. Our database consists of security log data, and the volume of the database is much bigger than previous works [53, 57].
@@ D.1 Graphs and Reports
3:$D\leftarrow$ empty map $\triangleright$ keys: path length, values: entity lists
@@
6:     $D[d]\leftarrow D[d]\cup\{e\}$triangleright bucket entity by distance
7:$d_{\max}\leftarrow\max\{d\mid d\in D\}$triangleright greatest distance observed
8:return $D[d_{\max}]$triangleright all entities farthest from $a_{e}$
@@ D.3 Split Train Test Set
We want to split the train and test set so that they have fewer overlaps. For example, with a simple graph $\{A-B-C,D-B-E\}$, it is best to split the paths in the subgraph $A-B-C$ and $D-B-E$ into two different sets, instead of $A-B$, $D-E$ to train, $A-C,D-E$ to test. For this purpose, we randomly split the training or test set and compute a customized total overlap score between every two paths from the train and test sets. We run this for T trials and select the split with the lowest overlap score.
@@
Thus, $S$ ranges from $-\beta$ (complete mismatch) to $\alpha$ (identical edge sets), and is exactly 00 when the two paths have no common edges at all.
@@
Table 5: Results with Master-Slave testing and finetuning. Models related to the MM testing is also included for comparison. denotes a special mixture of model use (MM).The average per incident and the total average reward are shown (Total average reward = sum of reward / total question count). We only should the related models in the table (Full result in Table 2). indicates the agent is instructed with additional format notes. indicates the agent is instructed with additional format notes. For finetune (Appendix E.2), incident 34 and 38 are the hold-out (test) incidents, and the performance drops significantly.
@@ E.3 Additional Experiment on Scaling
To further explore the limits of our test-time scaling method, we apply Best-of-N sampling to the baseline GPT-4o agent over 10 independent trials (see Figure 9). Across these trials, the mean reward rose from approximately 0.29 to 0.62. Although the reward has not fully converged by trial 10, the rate of improvement diminishes, indicating a flattening slope. Meanwhile, the average number of rounds continues to grow steadily. This follows that all our test-time in-context learning methods perform worse than baseline models in pass@k at high k values. A notion very similarly observed for RLVR approaches in [58].

### v2

@@ C.1 Features
Real-world security database Although the tenant is for a fictional company, the log data we use is from real-world tenant environments. Our database consists of security log data, and the volume of the database is much bigger than previous works Yang et al. [2023], Yu et al. [2018].
@@ D.1 Graphs and Reports
[…]
@@
6:  $D[d]\leftarrow D[d]\cup\{e\}$ $\triangleright$ bucket entity by distance
[…]
8:return $D[d_{\max}]$ $\triangleright$ all entities farthest from $a_{e}$
@@ D.3 Split Train Test Set
We want to split the train and test set so that they have fewer overlaps. For example, with a simple graph $\{A-B-C,D-B-E\}$, it is best to split the paths in the subgraph ABCA-B-C and DBED-B-E into two different sets, instead of ABA-B, DED-E to train, AC,DEA-C,D-E to test. For this purpose, we randomly split the training or test set and compute a customized total overlap score between every two paths from the train and test sets. We run this for T trials and select the split with the lowest overlap score.
@@
Thus, $S$ ranges from $\beta$ (complete mismatch) to $\alpha$ (identical edge sets), and is exactly 0 when the two paths have no common edges at all.
@@
Table 5: Results with Master-Slave testing and finetuning. Models related to the $M$ testing is also included for comparison. denotes a special mixture of model use ($M$).The average per incident and the total average reward are shown (Total average reward = sum of reward / total question count). We only should the related models in the table (Full result in Table 2). indicates the agent is instructed with additional format notes. indicates the agent is instructed with additional format notes. For finetune (Appendix E.2), incident 34 and 38 are the hold-out (test) incidents, and the performance drops significantly.
@@ E.3 Additional Experiment on Scaling
To further explore the limits of our test-time scaling method, we apply Best-of-N sampling to the baseline GPT-4o agent over 10 independent trials (see Figure 9). Across these trials, the mean reward rose from approximately 0.29 to 0.62. Although the reward has not fully converged by trial 10, the rate of improvement diminishes, indicating a flattening slope. Meanwhile, the average number of rounds continues to grow steadily. This follows that all our test-time in-context learning methods perform worse than baseline models in pass@k at high k values. A notion very similarly observed for RLVR approaches in [Yue et al., 2025].